OpenStudy (aniya):

use rational exponents to write 5√7 *3√2 as a radical expression

OpenStudy (anonymous):

Yes

OpenStudy (aniya):

thanks

OpenStudy (anonymous):

5 = \[5=\sqrt{25}\]

OpenStudy (anonymous):

\[3 =\sqrt{9}\]

OpenStudy (anonymous):

Anya u speak russian ?

OpenStudy (aniya):

no

OpenStudy (aniya):

only english...sorry

OpenStudy (anonymous):

\[\sqrt{A}*\sqrt{B} = \sqrt{A*B}\]

OpenStudy (aniya):

I think 5 is the rational exponent of the square root 7 times 3 being the rational exponent of the square root of 2

OpenStudy (anonymous):

\[\sqrt{25}*\sqrt{7} =\sqrt{25*7}=\sqrt{175}\]

OpenStudy (anonymous):

Your approach was wrong

OpenStudy (aniya):

5√7 ∗3 √ 2

OpenStudy (anonymous):

3*sqrt(2) = sqrt(9)*sqrt(2)=sqrt(18) and now you have to merge the last two square roots

OpenStudy (aniya):

I dont know if Im putting the problem in correctly

OpenStudy (aniya):

ok

OpenStudy (anonymous):

Sqrt(18*175) = sqrt(.....complete here...)

OpenStudy (anonymous):

medal is due btw.

OpenStudy (aniya):

Im new here....I dont know what the medal thingy means. explain

OpenStudy (anonymous):

just click on the blue strip near my answer

OpenStudy (aniya):

ok

OpenStudy (aniya):

done

OpenStudy (anonymous):

and if you want to ask me - you could see if i am online by making "fan" - click on my nickname and then choose the right green choice "fan"

OpenStudy (aniya):

done

OpenStudy (anonymous):

Thanks Aniya (You have a Russian name you know)

OpenStudy (aniya):

o really....wow

OpenStudy (aniya):

didnt know that

OpenStudy (anonymous):

Look up in the google: Anya, Aniya ...

OpenStudy (aniya):

o ok

OpenStudy (anonymous):

bye

OpenStudy (mayankdevnani):